# Attention-Guided Adaptation for Code-Switching Speech Recognition

### 14 Dec 2023 v1

@@ 1 Introduction
It is common for individuals to use multiple languages for verbal communication in their daily lives. A challenging issue in automatic speech recognition (ASR) is to handle the seamless interchange over languages in a speech utterance, known as the code-switching speech [1], where the semantics from human may be precisely expressed. The code-switching speech can be inter-sentential or intra-sentential [2], whereas the intra-sentential case is especially prevalent when a single sentence contains more than one language. To achieve the desirable performance in ASR, it is crucial to cope with the code-switching inputs in an encoder-decoder framework. In recent years, there has been a rapid advance in the development of foundation models for multilingual ASR such as XLS-R [3], Whisper [4], USM [5], and MMS [6], which have been pre-trained from large corpora of audio data with numerous languages. However, the capability of handling code-switching inputs using these backbone models remains unclear despite their proficiency in capturing inter-sentential speech sentences from multiple languages. In [7], the limited performance of multilingual language models in code-switching text data was shown. The code-switching speech recognition using multilingual speech models should be also bounded if code-switching scheme is not carefully designed.
@@
Recently, code-switching ASR has obtained significant attention, and can be categorized into two ways through either utilizing a shared encoder to create similar representations or adopting the separate encoders to establish distinct language embeddings. In [8, 9], a shared encoder was incorporated with language identification (LID) in a training procedure. In [10], the similarity of token embeddings in a monolingual language was explicitly encouraged to be similar. In addition, the shared encoder [11] was extended to develop independent self-attention networks for individual languages. However, in [12, 13], the separate encoders were exploited by utilizing transformer-transducer and CTC model for different languages. While separate encoders improved the recognition results, the additional memory and computation were required.
@@
This paper develops a new approach to code-switching ASR by adapting the pre-trained multilingual ASR model using Whisper based on adapters. A parameter-efficient learning based on an attention-guided adaptation is proposed. Initially, a behavior analysis of the attention maps in the transformer decoder is conducted. A head selection method is accordingly proposed to identify those LID token attention heads. These selected heads are then adapted by minimizing an attention-guided loss to ensure the corresponding LID tokens to be correctly attended. In addition, a two-stage adapter training process is designed to carry out a stable parameter-efficient learning (PEL). The experiments on code-switching Mandarin and English speech recognition show the efficiency and effectiveness using the proposed method.
@@ 2 Code-Switching in Whisper
To investigate how LID tokens behave in a bilingual code-switching ASR task, this study first conducts an insightful analysis on the attention map of each head in a decoder layer. It is interesting to find that there are three typical patterns including self attention, neighboring attention and special token attention [15] as shown by the examples using Whisper model in Figure 2 (left and middle subfigures). Notably, a distinctive pattern called the LID token attention is explored as shown in the right subfigure. This paper is accordingly motivated to guide the model learning through minimizing an attention-guided loss with the goal of highlighting the attention behaviors in code-switching speech so that the bilingual ASR decoder is allowed to sensitively pay attention on the language changes in speech signal for word prediction.

### 12 Jan 2024 v2

@@ 1 Introduction
It is common for individuals to use multiple languages for verbal communication in their daily lives [1]. A challenging issue in automatic speech recognition (ASR) is to handle the seamless interchange over languages in a speech utterance, known as the code-switching speech [2], where the semantics from human may be precisely expressed. The code-switching speech can be inter-sentential or intra-sentential [3], whereas the intra-sentential case is especially prevalent when a single sentence contains more than one language. To achieve the desirable performance in ASR, it is crucial to cope with the code-switching inputs in an encoder-decoder framework. In recent years, there has been a rapid advance in the development of foundation models for multilingual ASR such as XLS-R [4], Whisper [5], USM [6], and MMS [7], which have been pre-trained from large corpora of audio data with numerous languages. However, the capability of handling code-switching inputs using these backbone models remains unclear despite their proficiency in capturing inter-sentential speech sentences from multiple languages. In [8], the limited performance of multilingual language models in code-switching text data was shown. The code-switching speech recognition using multilingual speech models should be also bounded if code-switching scheme is not carefully designed.
@@
Recently, code-switching ASR [9] has obtained significant attention, and can be categorized into two ways through either utilizing a shared encoder to create similar representations or adopting the separate encoders to establish distinct language embeddings. In [10, 11], a shared encoder was incorporated with language identification (LID) in a training procedure. In [12, 13], the similarity of token embeddings in a monolingual language was explicitly encouraged to be similar. In addition, the shared encoder [14, 15, 16] was extended to develop independent self-attention networks for individual languages. However, in [17, 18, 19], the separate encoders were exploited by utilizing transformer-transducer and CTC model for different languages. While separate encoders improved the recognition results, the additional memory and computation were required.
@@
This paper develops a new approach to code-switching ASR by adapting the pre-trained multilingual ASR model using Whisper based on adapters. A parameter-efficient learning based on an attention-guided adaptation is proposed. Initially, a behavior analysis of the attention maps in the transformer decoder is conducted. A head selection method is accordingly proposed to identify those LID token attention heads. These selected heads are then adapted by minimizing an attention-guided loss to ensure the corresponding LID tokens to be correctly attended. In addition, a two-stage adapter training process is designed to carry out a stable parameter-efficient learning [20]. The experiments on code-switching Mandarin and English speech recognition show the efficiency and effectiveness using the proposed method.
@@ 2 Code-Switching in Whisper
To investigate how LID tokens behave in a bilingual code-switching ASR task, this study first conducts an insightful analysis on the attention map of each head in a decoder layer. It is interesting to find that there are three typical patterns including self attention, neighboring attention and special token attention [22, 23] as shown by the examples using Whisper model in Figure 2 (left and middle subfigures). Notably, a distinctive pattern called the LID token attention is explored as shown in the right subfigure. This paper is accordingly motivated to guide the model learning through minimizing an attention-guided loss with the goal of highlighting the attention behaviors in code-switching speech so that the bilingual ASR decoder is allowed to sensitively pay attention on the language changes in speech signal for word prediction.